# version 0.6

### v3

@@ 13.1 Spectral filtering in One Dimension
$=\beta\cdot\mu_{\alpha}(\mathbf{u}_{t-1:1}),$ (13.1)
@@
$\mu_{\alpha}(\mathbf{u}_{t-1:1})=\mathbf{u}_{t-1}+\sum_{i=1:t-1}(\alpha-1)\alpha^{i-1}\mathbf{u}_{t-1-i}=\mu_{\alpha}^{\top}\tilde{\mathbf{u}}_{t-1}.$ (13.2)
@@
$\mu_{\alpha}=(\alpha-1)\begin{bmatrix}\frac{1}{\alpha-1}&1&\alpha&\alpha^{2}&\ldots&\alpha^{T-1}\end{bmatrix}\in{\mathbb{R}}^{T},$
@@
$\tilde{\mathbf{u}}_{t}=\begin{bmatrix}\mathbf{u}_{t}&\ldots&\mathbf{u}_{1}&0\ldots&0\end{bmatrix}\in{\mathbb{R}}^{T}.$
@@ 13.2 Spectral Predictors
Define the class of spectral linear predictors of order $k$ as
$\prod_{h}^{\lambda}=\left\{\pi_{M}\ \ \left|\ \ \pi_{M}(\mathbf{u}_{1:t},\mathbf{y}_{1:t})=\sum_{i=1}^{h}M_{i}\cdot\phi_{i}(\tilde{\mathbf{u}}_{t-i})\right.\right\}.$
Here the operator $\phi_{i}(\mathbf{u}_{t-1})$ takes the inner product of the $i$’th eigenvector of $Z_{T}$ with $\tilde{\mathbf{u}}_{t-1}$, where $\tilde{\mathbf{u}}_{t}=\begin{bmatrix}\mathbf{u}_{t}&\ldots&\mathbf{u}_{1}&0\ldots&0\end{bmatrix}\in{\mathbb{R}}^{T}$.
@@
Let $\{\phi_{j}\}$ be the eigenvalues of Z. Then for all $j\in[T]$, ,$\alpha\in[0,1]$, we have
$|\phi_{j}^{\top}\mu_{\alpha}|=O\left(T^{2}e^{-\frac{j}{\log T}}\right).$
@@
$\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t-1}^{\pi_{1}}$ $=\beta\cdot\mu_{\alpha}(\mathbf{u}_{t-1:1})$
@@
$M_{i}^{\star}=\beta\phi_{i}^{\top}\mu_{\alpha}=O(T^{2}e^{-\frac{i}{\log T}}).$
@@
$\leq\sum_{i=h+1}^{T}\|M_{i}^{\star}\|\|\phi_{i}\|_{1}\|\tilde{\mathbf{u}}_{t-1}\|_{\infty}$ Holder inequality
@@
Observe that
$\|\mu_{\alpha}\|_{2}^{2}=\sum_{t=0}^{T-1}(1-\alpha)^{2}\alpha^{2i}\leq\frac{(1-\alpha)^{2}}{1-\alpha^{2}}\leq 1-\alpha$
implying that $g(\alpha)=(\mu_{\alpha}^{\top}\phi_{j})^{2}\leq\|\phi_{j}\|^{2}\|\mu_{\alpha}\|^{2}\leq 1-\alpha$. Similarly, we have
@@
$=8T^{2}\sigma_{j}$
$\leq O(T^{2}e^{-\frac{j}{\log T}})$ Theorem 13.1
@@ 13.2.1 Generalization to High Dimension
$\hat{\mathbf{y}}^{\pi}_{t}-\mathbf{y}_{t-1}-CB\mathbf{u}_{t-1}$
@@
$=\sum_{i=1}^{t-1}CU^{\top}(\sum_{j=1}^{d_{x}}(\alpha^{i-1}_{j}-\alpha_{j}^{i-2})e_{i}e_{i}^{\top})UB\mathbf{u}_{t-i-1}$
$=\sum_{j=1}^{d_{x}}\beta_{j}\sum_{i=1}^{t-1}(\alpha^{i-1}_{j}-\alpha_{j}^{i-2})\mathbf{u}_{t-1-i}$ for scalars $\beta_{j}$
$=\sum_{j=1}^{d_{x}}{M}^{\star}_{j}\mu_{\alpha_{j}}(\mathbf{u}_{1:t-1})$
Thus, we can apply the same approximation techniques from the previous section, and derive an analoguous conclusion to Lemma 13.3 for dimension higher than one. The details are left as exercise 13.5.
@@ 13.3 Online Spectral Filtering
The prediction class $\prod_{h}^{\lambda}$ can be learned by online gradient descent in the online convex optimization framework, giving rise to the following algorithm.
@@
4:     Predict $\hat{\mathbf{y}}_{t}=\sum_{j=1}^{h}M^{t}_{j}\phi_{j}(\tilde{\mathbf{u}}_{t-i})$ , as per definition 13.2, and observe true $\mathbf{y}_{t}$.
@@
The online gradient descent theorem, as per the previous chapter, directly gives the following corollary,
@@
For choice of step sizes $\eta_{t}=O(\frac{1}{\sqrt{t}})$, we have that
$\sum_{t}\|\mathbf{y}_{t}-\hat{\mathbf{y}}_{t}\|^{2}-\min_{\pi\in\prod_{h}^{\lambda}}\|\mathbf{y}_{t}-\mathbf{y}_{t}^{\pi}\|^{2}=O(\sqrt{T}).$
@@ 13.4 Bibliographic Remarks
Under probabilistic assumptions, the work of [HMR16] shows that the matrices $A,B,C,D$ of an unknown and partially observed time-invariant linear dynamical system can be learned using first-order optimization methods.
@@
For an extension of the spectral learning technique to asymmetric matrices, see [HLS+18].
@@ 13.5 Exercises
Appendix A Concepts from Online Convex Optimization
In this section we give a brief introduction to the theory of Online Convex Optimization. The reader is referred to [Haz16] for a comprehensive survey.
$\mathbf{x}_{t}^{\mathcal{A}}={\mathcal{A}}(f_{1},...,f_{t-1})\in\mathcal{K}.$
We define the regret of ${\mathcal{A}}$ after $T$ iterations as:
$\mathrm{{regret}}_{T}({\mathcal{A}})=\sup_{\{f_{1},...,f_{T}\}\subseteq{\mathcal{F}}}\left\{\sum_{t=1}^{T}f_{t}(\mathbf{x}_{t}^{\mathcal{A}})-\min_{\mathbf{x}\in\mathcal{K}}\sum_{t=1}^{T}f_{t}(\mathbf{x})\right\}.$ (A.1)
A.1 Online Gradient Descent
Perhaps the simplest algorithm that applies to the most general setting of online convex optimization is online gradient descent. This algorithm, which is based on standard gradient descent from offline optimization. Pseudo-code for the algorithm is given in Algorithm 8.
Algorithm 8 online gradient descent
2:  for $t=1$ to $T$ do
3:     Play $\mathbf{x}_{t}$ and observe cost $f_{t}(\mathbf{x}_{t})$.
4:     Update and project:
$\mathbf{y}_{t+1}=\mathbf{x}_{t}-\eta_{t}\nabla f_{t}(\mathbf{x}_{t})$
$\mathbf{x}_{t+1}=\operatorname*{\arg\min}_{\mathbf{x}\in\mathcal{K}}\|\mathbf{x}-\mathbf{y}_{t+1}\|^{2}$
5:  end for
  • $D$ = Euclidean diameter of the decision set, i.e.
    $\max_{\mathbf{x},\mathbf{y}\in\mathcal{K}}\|\mathbf{x}-\mathbf{y}\|.$
    $\forall t\ ,\ \|\nabla f_{t}(\mathbf{x}_{t})\|\leq G.$
Theorem A.1.

### v4

@@ 11.1 Spectral filtering in One Dimension
$=\beta\mathbf{u}_{t-1}+\beta\cdot\mu_{\alpha}(\mathbf{u}_{t-1:1}),$ (11.1)
@@
$\mu_{\alpha}(\mathbf{u}_{t-1:1})=\sum_{i=1:t-1}(\alpha-1)\alpha^{i-1}\mathbf{u}_{t-1-i}=\mu_{\alpha}^{\top}\tilde{\mathbf{u}}_{t-1}.$ (11.2)
@@
$\mu_{\alpha}=(\alpha-1)\begin{bmatrix}1&\alpha&\alpha^{2}&\ldots&\alpha^{T-1}\end{bmatrix}\in{\mathbb{R}}^{T},$
@@
$\tilde{\mathbf{u}}_{t}=\begin{bmatrix}\mathbf{u}_{t-1}&\ldots&\mathbf{u}_{1}&0\ldots&0\end{bmatrix}\in{\mathbb{R}}^{T}.$
@@ 11.2 Spectral Predictors
Define the class of spectral linear predictors of order $k$ recursively as
$\prod_{h}^{\lambda}=\left\{\pi_{M}\ \ \left|\ \ \pi_{M}(\mathbf{u}_{1:t},\mathbf{y}_{1:t})-\pi_{M}(\mathbf{u}_{1:t-1},\mathbf{y}_{1:t-1})=M_{0}\mathbf{u}_{t}+\sum_{i=1}^{h}M_{i}\cdot\phi_{i}(\tilde{\mathbf{u}}_{t-1})\right.\right\}.$
Here the operator $\phi_{i}(\mathbf{u}_{t-1})$ takes the inner product of the $i$’th eigenvector of $Z_{T}$ with $\tilde{\mathbf{u}}_{t-1}$, where $\tilde{\mathbf{u}}_{t}=\begin{bmatrix}\mathbf{u}_{t-1}&\ldots&\mathbf{u}_{1}&0\ldots&0\end{bmatrix}\in{\mathbb{R}}^{T}$.
@@
Let $\{\phi_{j}\}$ be the eigenvectors of Z. Then there exists constants $c_{1},c_{2}>0$ such that for all $j\in[T]$, $\alpha\in[0,1]$, we have
$|\phi_{j}^{\top}\mu_{\alpha}|\leq c_{1}e^{-\frac{c_{2}j}{\log T}}.$
@@
$\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t-1}^{\pi_{1}}-\beta\mathbf{u}_{t}$ $=\beta\cdot\mu_{\alpha}(\mathbf{u}_{t-1:1})$
@@
$M_{i}^{\star}=\beta\phi_{i}^{\top}\mu_{\alpha}=O(e^{-\frac{i}{\log T}}).$
@@
$\leq\sum_{i=h+1}^{T}\|M_{i}^{\star}\|\|\phi_{i}\|_{2}\|\tilde{\mathbf{u}}_{t-1}\|_{2}$ Cauchy-Schwartz
@@
Consider the scalar function $g(\alpha)=(\mu_{\alpha}^{\top}\phi_{j})^{2}$ over the interval $[0,1]$. First, notice that by definition of $\phi_{j}$ as the eigenvectors of $Z_{T}$, and $\sigma_{j}$ as the corresponding eigenvalues, we have
[…]
On the other hand, we claim that the function $g(\alpha)$ is L-Lipschitz for a constant value of $L$. To see this, first notice that
$\|\mu_{\alpha}\|_{2}^{2}=\sum_{i=0}^{T-1}(1-\alpha)^{2}\alpha^{2i}\leq\frac{(1-\alpha)^{2}}{1-\alpha^{2}}=\frac{1-\alpha}{1+\alpha}\leq 1-\alpha,$
implying that $g(\alpha)=(\mu_{\alpha}^{\top}\phi_{j})^{2}\leq\|\phi_{j}\|^{2}\|\mu_{\alpha}\|^{2}\leq 1-\alpha$. Next, we have
@@
$\leq 2\|\mu_{\alpha}\|_{2}\|\phi_{j}\|_{2}\left\lvert\frac{\partial}{\partial{\alpha}}(\mu_{\alpha}^{\top}\phi_{j})\right\rvert$ Cauchy-Schwarz
@@
$=12\sigma_{j}$
$\leq 12c_{1}e^{-\frac{c_{2}j}{\log T}}$ Theorem 11.1
@@ 11.2.1 Generalization to High Dimension
$\mathbf{y}_{t}-\mathbf{y}_{t-1}-CB\mathbf{u}_{t-1}$
@@
$=\sum_{i=1}^{t-1}CU^{\top}(\sum_{j=1}^{d_{x}}(\alpha^{i-1}_{j}-\alpha_{j}^{i-2})e_{j}e_{j}^{\top})UB\mathbf{u}_{t-i-1}$
$=\sum_{j=1}^{d_{x}}c_{j}b_{j}^{\top}\sum_{i=1}^{t-1}(\alpha^{i-1}_{j}-\alpha_{j}^{i-2})\mathbf{u}_{t-1-i}$ $c_{j},b_{j}\leftarrow j^{th}\mbox{ row/col of }CU^{\top},UB$
$=\sum_{j=0}^{d_{x}}\left(M^{\star}_{j}\mathbf{u}_{t}+{M}^{\star}_{j}\mu_{\alpha_{j}}(\mathbf{u}_{1:t-1})\right)$
$=\tilde{M}^{\star}\mathbf{u}_{t}+\tilde{M}^{\star}\mu_{\alpha_{j}}(\mathbf{u}_{1:t-1})$
Thus, we can apply the same approximation techniques from the previous section, and derive an analogous conclusion to Lemma 11.3 for dimension higher than one. Crucially, the number of parameters required does not depend on the hidden dimension of the original system. The details are left as exercise 11.6.
@@ 11.3 Online Spectral Filtering
The prediction class $\prod_{h}^{\lambda}$ can be learned by online gradient descent in the online convex optimization framework, as per Algorithm 3, giving rise to the following algorithm.
@@
4:     Predict $\hat{\mathbf{y}}_{t}=M_{0}^{t}\mathbf{u}_{t}+\sum_{j=1}^{h}M^{t}_{j}\phi_{j}(\tilde{\mathbf{u}}_{t-i})$ , as per definition 11.2, and observe true $\mathbf{y}_{t}$.
@@
The regret guarantees of the online gradient descent algorithm, as given in Theorem 6.1, directly implies the following corollary,
@@
Let $D$ be the diameter of the parameter set $\mathcal{K}$, and $G$ an upper bound on the norm of the gradients for the loss functions $f_{t}$ as in Algorithm 7. For choice of step sizes $\eta_{t}=\frac{D}{G\sqrt{t}}$, we have that
$\sum_{t}\|\mathbf{y}_{t}-\hat{\mathbf{y}}_{t}\|^{2}-\min_{\pi\in\prod_{h}^{\lambda}}\|\mathbf{y}_{t}-\mathbf{y}_{t}^{\pi}\|^{2}\leq 2GD\sqrt{T}.$
11.4 Conclusion
In this chapter, we introduced Spectral Filtering as a novel approach to state estimation and prediction in linear dynamical systems, offering an alternative to classical methods like Kalman filtering and online learning of linear predictors. By leveraging spectral decompositions, we bypass the need for explicit system identification and provide a framework that is robust to system stability properties.
We began by outlining the core challenges of traditional state estimation approaches. Kalman filtering, while optimal under Gaussian noise assumptions, requires knowledge of system matrices, which may be difficult to estimate in high-noise or adversarial settings. Online learning of linear predictors relaxed these requirements but introduced a dependence on system stability constants, which can be large and unknown. Spectral Filtering circumvents these limitations by constructing predictors in a transformed basis, allowing for efficient representation without requiring explicit system matrices.
We demonstrated the method first in the one-dimensional case, providing intuition for how spectral predictors extract relevant dynamical structure. This was then extended to higher-dimensional settings, where spectral filtering enables a structured and computationally efficient approach to forecasting. Crucially, we showed that this approach does not rely on system stability constants, making it particularly appealing for learning in both stable and unstable dynamical systems.
From a methodological standpoint, we connected Spectral Filtering to convex optimization techniques, showing that it can be implemented via online learning in a spectral basis. This formulation allows for efficient sequential learning and prediction, with provable performance guarantees.
Looking ahead, Spectral Filtering opens the door for several exciting directions. Extensions to asymmetric system matrices, as referenced in the bibliographic section, provide a pathway toward broader applicability. Moreover, integrating spectral methods with control strategies may lead to novel approaches for nonstochastic control with partial observations, which we explore in the next chapter.
@@ 11.5 Bibliographic Remarks
Under probabilistic assumptions, the work of [HMR18] shows that the matrices $A,B,C,D$ of an unknown and partially observed time-invariant linear dynamical system can be learned using first-order optimization methods.
@@
The first extension of the spectral learning technique to asymmetric matrices is due to [HLS+18]. A more recent advancement of spectral filtering to general linear dynamical systems was recently put forth in [MH25].
The technique of spectral filtering has found applications in control and sequence prediction since its introduction.
In [AHL+18], the authors leverage a spectral filtering technique to represent linear dynamical systems concisely in a linear basis, formulating the optimal control problem as a convex program. This method circumvents the non-convex challenges associated with explicit system identification and latent state estimation, providing provable optimality guarantees for the control signal.
A recent line of work makes use of the spectral basis discovered in the line of work surveyed in this chapter in order to design neural architectures for sequence prediction. This was first suggested in the context of state space models in the spectral basis in [ASCH23], and extended to hybrid models [LND+24]. Later extensions explored efficient implementation [ACD+24] and the concept of length generalization in sequence prediction [MDA+24].
@@ 11.6 Exercises
11.3. The fact that $Z_{T}$ is positive semidefinite is important to having the eigenvalues of $Z_{T}$ decay. Give an example of a family of $n\times n$ non-positive semidefinite Hankel matrices that have eigenvalues with constant absolute value.
11.4. The view of Theorem 13.5 is a bit misleading, as it compares the iterates of Online Gradient Descent with the hypothesis class $\Pi_{h}^{\lambda}$ and uses big-Oh notation to obscure the diameter and Lipschitz constants of the predictors. Assume that $d_{x}=1$.
  1. (i)
    Find a convex set $\mathcal{M}$ of tuples $(M_{1},M_{2},\dots)$’s to optimize across such that its diameter is some $\textrm{polylog}(T,1/\varepsilon)$. The norm we measure diameter with respect to is the sum of the Frobenius norms of the $M_{i}$’s.
  2. (ii)
    One can further show that the norm of the gradient is at most $\textrm{polylog}(T,1/\varepsilon)$ as well. Use these two facts to show that with the right value of $\varepsilon$, the regret against the hypothesis class $\Pi_{\star}$ is at most $\tilde{O}(\sqrt{T})$.
Part IV Online Control with Partial Observation
Chapter 12 Policy Classes for Partially Observed Systems
In previous chapters, we explored control and learning techniques in fully observed linear dynamical systems, where the system state is directly accessible to the controller. However, in many real-world scenarios—such as robotics, economics, and signal processing—the state is only partially observed, requiring the controller to infer system dynamics from noisy or incomplete observations.
This chapter focuses on partially observed LDS (PO-LDS), where the state is hidden, and only a linear projection of it is observed. For the remainder of this part of the text, we restrict ourselves to time-invariant systems, and refer to the bibliographic section for extensions. The system follows the general form:
$\mathbf{x}_{t+1}$ $=A\mathbf{x}_{t}+B\mathbf{u}_{t}+\mathbf{w}_{t}$
$\mathbf{y}_{t}$ $=C\mathbf{x}_{t}.$
Here, $\mathbf{x}_{t}$ represents the hidden system state, $\mathbf{u}_{t}$ the control input, $\mathbf{y}_{t}$ the observed output, and $\mathbf{w}_{t}$ the disturbance. Often, partially observed LDSs are defined with an additional noise term corrupting the observations. Exercise 12.6 asks the reader to prove that even in such circumstances we may assume the previously stated form for partially observed systems without loss of any generality.
Unlike in the fully observed setting, where controllers can react based on the full state, partial observability introduces fundamental limitations on policy design.
Not all control policies designed for fully observed systems extend naturally to partially observed settings. In this chapter, we systematically explore different policy classes, analyzing their expressivity, computational tractability, and suitability for PO-LDS.
  • Linear Observation-Action Controllers (LOACs): These controllers directly map observations to actions but are highly inexpressive, as they ignore past observations and fail to capture hidden state dependencies.
  • Generalized Linear Observation-Action Controllers (GOACs): These extend LOACs by incorporating past history of observations, significantly improving expressiveness.
  • Linear Dynamic Controllers (LDCs): These policies model internal system dynamics, making them more powerful than simple observation-based policies.
  • Disturbance Response Controllers (DRCs): Introduced as an alternative to state-based policies, DRCs directly act on a counterfactual signal—the “Nature’s y’s”—which captures the system’s response had no control been applied. We show that DRCs can approximate LDCs, making them a powerful convex relaxation.
We conclude by analyzing the relationships between these policy classes, showing how they fit into the broader framework of nonstochastic control with partial observation. This chapter lays the groundwork for online learning and regret minimization in PO-LDS, which we explore in the next chapter.
12.1 Linear Observation Action Controllers
Definition 12.1.
A linear observation-action controller $K\in{\mathbb{R}}^{d_{u}\times d_{y}}$ is a linear operator mapping observation to control as $\mathbf{u}_{t}^{\pi}=K\mathbf{y}_{t}$. We denote by $\Pi_{\gamma}^{L}$ the set of all linear controllers bounded in Fronbenius norm by $\|K\|\leq\gamma$, that are also $\gamma$-stabilizing for given LDS.
While this class of policies is natural to consider, it turns out to be very inexpressive. Even for optimal control with stochastic noise, the optimal policy cannot be approximated by a linear observation-action policy (see Exercise 12.6). Therefore we proceed to define more expressive and useful policies. However, LOACs will be useful in simplifying proofs later on.
12.1.1 Generalized Linear Observation Action Controllers
We can generalize linear observation-action controllers to apply to a history of past observations. This additional generality, of taking a history of observations, greatly increases the expressive power of this policy class. The formal definition is below.
Definition 12.2.
Generalized linear observation-action controllers parameterized by matrices $K_{0:h}$ choose a control input as
$\mathbf{u}_{t}^{\pi}=\sum_{i=0}^{h}K_{i}\mathbf{y}_{t-i}.$
Denote by $\Pi^{G}_{h,\gamma}$ be that class of all such policies satisfying $\sum_{i=0}^{h}\|K_{i}\|\leq\gamma$ that are also $\gamma$-stabilzing.
Generalized observation-action policies can be seen as a special case of observation-action controllers using the technique of dimension lifting as follows. Consider any partially-observed system $(A,B,C)$ and perturbation sequence $\mathbf{w}_{t}$. Consider the lifted system $(\tilde{A},\tilde{B},\tilde{C},\tilde{D})$ given by
$\mathbf{z}_{t+1}=\tilde{A}\mathbf{z}_{t}+\tilde{B}\mathbf{u}_{t}+\tilde{D}\mathbf{w}_{t},\quad\tilde{\mathbf{y}}_{t}=C\mathbf{z}_{t},$
where
$\tilde{A}=\begin{array}[]{|c|c|c|c|c|}\hline\cr A&0&\ldots&0&0\\ \hline\cr I&0&\ldots&0&0\\ \hline\cr 0&I&\ldots&0&0\\ \hline\cr\vdots&\vdots&\cdots&\vdots&\vdots\\ \hline\cr 0&0&\cdots&I&0\\ \hline\cr\end{array}\ \ ,\ \ \tilde{B}=\begin{array}[]{|c|}\hline\cr B_{t}\\ \hline\cr 0\\ \hline\cr 0\\ \hline\cr\vdots\\ \hline\cr 0\\ \hline\cr\end{array}\ \ ,$
$\ \ \tilde{C}=\begin{array}[]{|c|c|c|c|}\hline\cr C&C&\dots&C\\ \hline\cr\end{array}\ \ ,\ \ \tilde{D}=\begin{array}[]{|c|}\hline\cr I\\ \hline\cr 0\\ \hline\cr 0\\ \hline\cr\vdots\\ \hline\cr 0\\ \hline\cr\end{array}.$
The following lemma relates the original and lifted system (for a proof see exercise 12.6).
Lemma 12.3.
12.2 Linear Dynamic Controllers
The most well-known class of controllers for partially observable linear dynamical systems is that of linear dynamical controllers due to its connection to Kalman filtering.
Note that LDCs are $\varepsilon$-approximated by generalized observation-action controllers, namely GOACs. The proof is left as Exercise 12.6.
Lemma 12.5.
Consider a partially-observed time-invariant linear dynamical system $\{A,B,C\}$. The class $\Pi^{G}_{h^{\prime},\gamma^{\prime}}$ $\varepsilon$-approximates the class $\Pi^{LDC}_{\gamma}$ for $h^{\prime}=O\left(\gamma\log\frac{\gamma}{\varepsilon}\right),\gamma^{\prime}=2\gamma^{3}$.
12.3 Disturbance Response Controllers
The most important policy class in the context of nonstochastic control is that of Disturbance Response Controllers, or DRC. These policies act on a signal called Nature’s y’s, denoted $\mathbf{y}^{\mathbf{nat}}_{t}$, which is the would-be observation at time $t$ assuming all inputs to the system were zero from the beginning of time.
To simplify presentation, we only consider linear time invariant linear dynamical systems that are stable. This allows us to ignore the stabilizing controller component of disturbance action control and greatly simplifies the exposition of the main ideas. Extension to time-varying systems is explored in the exercises.
Definition 12.6 (Disturbance Response Controller).
A disturbance response policy for stable systems $\pi_{M_{1:h}}$ is parameterized by the matrices $M_{0:h}=[{M}_{0},\ldots,{M}_{h}]$. It outputs control $\mathbf{u}_{t}^{\pi}$ at time $t$ according to the rule
$\mathbf{u}_{t}^{\pi}=\sum_{i=0}^{h}{M}_{i}\mathbf{y}^{\mathbf{nat}}_{t-i},$
$\mathbf{x}^{\mathbf{nat}}_{t+1}$ $=A\mathbf{x}^{\mathbf{nat}}_{t}+\mathbf{w}_{t}=\sum_{i=0}^{t}A^{i}\mathbf{w}_{t-i}\ ,\ \mathbf{y}^{\mathbf{nat}}_{t+1}=C\mathbf{x}^{\mathbf{nat}}_{t+1}.$
Denote by $\Pi^{R}_{h,\gamma}$ the set of all disturbance response policies that are $\gamma$-stabilizing 111for example, by having additional linear controllers $K_{1:T}$ that are $\gamma$-stabilizing linear controllers for the system and
$\sum_{i=0}^{h}\|M_{i}\|\leq\gamma.$
The main advantage of this policy class, as opposed to previously considered policies in optimal control, is that they yield convex parametrization. Thus, this policy class allows for provably efficient optimization and algorithms that we explore in the next chapter.
We conclude with the following important lemma on an expression that is important to compute nature’s y’s. The proof of this lemma follows directly from the definition and is left as exercise 12.6.
Lemma 12.7.
$\mathbf{y}^{\mathbf{nat}}_{t}=\mathbf{y}_{t}-C\sum_{i=1}^{t}A^{i}B\mathbf{u}_{t-i}$
$\mathbf{z}_{t+1}$ $=A\mathbf{z}_{t}+B\mathbf{u}_{t},$
$\mathbf{y}^{\mathbf{nat}}_{t}$ $=\mathbf{y}_{t}-C\mathbf{z}_{t}.$
12.3.1 Expressivity of Disturbance Response Controllers
A DRC always produces controls that ensure the system remains state-bounded. This can be shown analogously to Lemma 6.6 (see Exercise 12.6).
To demonstrate that DRCs approximate GOACs, we instead prove that LOACs, a weaker class of policies, are $\varepsilon$-approximated by DRCs. Nevertheless, this observation is sufficient to conclude the stronger claim, given that GOACs can be written as LOACs on a lifted system as noted in Lemma 12.3, and simultaneously that a DRC with history $H^{\prime}$ on the lifted system can be written as a DRC with history $H^{\prime}+h$ over the original system.
Lemma 12.8.
Consider a $\gamma$-stable time-invariant linear dynamical system $(A,B,C)$. The class $\Pi^{R}_{h^{\prime},\gamma^{\prime}}$ $\varepsilon$-approximates the class $\Pi^{L}_{\gamma}$ for parameters $h^{\prime}=O\left(\gamma\log\frac{\gamma}{\varepsilon}\right)$, $\gamma^{\prime}=O(\gamma^{6})$.